OpenStudy (deadshot):

What is the equation of the line between the points (2, 8) and (3, -2)?

hartnn (hartnn):

first find slope m

OpenStudy (anonymous):

(y-y1) =m(x-x1) where m is the slope = (y2-y1) / (x2-x1)

OpenStudy (deadshot):

so the slope is -10?

hartnn (hartnn):

yes,

hartnn (hartnn):

now you have a point and slope use y-y1 = m (x-x1) use x1,y1 as any one of the 2 points given

OpenStudy (deadshot):

so then it would be -2-8=-10(3-2) correct?

OpenStudy (uri):

Nope,  y-8=-10(x-2)